# Macroscopic States in Bose–Einstein Condensate Dark Matter Model with Axionlike Interaction

### v1

@@ Abstract
The phase diagrams of ultralight dark matter (DM), modeled as a self-gravitating Bose–Einstein condensate with axionlike interaction, are studied. We classify stable, metastable and unstable DM states over a wide range of condensate wave function amplitudes. It is shown that the axionlike interaction causes instability and an imaginary speed of sound at low amplitudes, whereas in a specific high-amplitude band, DM attains a stable state capable of forming a dense solitonic core and suppressing quantum fluctuations in the surrounding galactic DM halo. These findings are corroborated by evaluating thermodynamic functions for DM in the dwarf galaxy NGC 2366 and in its hypothetical analogues with different core-to-halo mass ratios. Distinct DM phase compositions respond differently to fluctuation-induced partial pressure, resulting in a first-order phase transition in a certain range of an interaction parameter. While the DM properties in NGC 2366 lie within the supercritical regime, the phase transition nonetheless provides a thermodynamic marker separating stable from unstable DM configurations. Once a dense core forms–reaching a threshold of about 12% of the total mass–the enhanced gravitation stabilizes the DM halo against fluctuations, while the internal pressure ensures core stability. In particular, we find that NGC 2366’s dense DM comprises roughly 19% of the DM mass while occupying only 4.7% of its total volume.
@@ I Introduction
One may observe parallels with the ensemble of oscillons Osc1 generated by the instanton self-interaction, leading in the absence of gravity to the three-dimensional sine-Gordon equation Man04 . Moreover, established results on axitons BZ19 ; KT94 ; ax5 are valuable for further comparison of sets of states.
Neglecting the Universe’s evolution in describing current visible structures, we thus focus on the statistical properties of self-interacting DM (SIDM) within the model of self-gravitating Bose–Einstein condensate (BEC). Phase diagrams should then reflect these properties, tracing transitions among distinct states and configurations.
Numerous studies have rigorously examined the properties of BEC DM with two- and three-particle interactions, using both exact numerical methods and various analytical approximations, such as the Thomas–Fermi approximation and the variational method (see Chavanis2 ; Peebles ; Bohmer ; Harko2011 ; Zang2018 ; Nazar ; SV24 ). The thermodynamic functions in these models have been carefully studied, and a new aspect of DM has emerged GKN20 ; GN21 ; GN23 : the prediction of a first-order phase transition between rarefied and dense states, which is driven by quantum fluctuation pressure and governed by coupling constants. Additionally, we would like to note the impact of the formation of composites from the initial DM particles GN23 ; GN22 ; GKKN , as well as the quantum entanglement of parts of the DM halo GN21 ; Lee18 .
@@
To constrain the parameters of our BEC DM model described by the Gross–Pitaevskii–Poisson equations, we consider the DM dominating dwarf galaxy NGC 2366 as a reference. This nearby galaxy provides an opportunity to study reionization processes in vigorous low-metallicity starbursts, such as those that occurred during the cosmic reionization epoch Iz99 . In this regard, we intend to calibrate the model by fitting its observed rotation curve Oh08 , and then generate a statistical ensemble of hypothetical DM distributions following from the model equations under similar conditions. By comparing different configurations using phase diagrams, we aim to reveal the uniqueness of the DM structure in NGC 2366, which forms the gravitational background for subgalactic processes.
The paper is structured as follows. In Sec. II we introduce the main equations of the model and, through various simplifications, analyze the properties essential for interpreting subsequent findings. In Sec. III we explore nonperturbative solutions and their application to modeling the DM distribution in the dwarf galaxy NGC 2366. We then present our main results – statistical characteristics and phase diagrams based on the numerically obtained DM profiles. The final Sec. IV contains a summary of the findings and concluding remarks.
@@ II.1 The model equations
It immediately leads to the dimensionless measure of gravitational interaction:
@@
With $v=\hbar^{3}c/(mf^{2}_{\text{a}})$ defined via the axion decay constant $f_{\mathrm{a}}$, it is expected that in models featuring ultralight particles, $f_{\text{a}}$ may fall within the range $10^{18}\,\text{eV}<f_{\text{a}}<10^{21}\,\text{eV}$ Chavanis2 .
@@
In our study, $\nu$ is treated as a free parameter, reflecting the arbitrariness of $\tilde{\mu}$.
@@ II.2 The particular cases
The parameters $\nu_{1}$-$\nu_{5}$ are (locally) critical for the curves in Fig. 1. Their values are established through conditions:
@@
Let us analyze the $p-\mu$ phase diagram in Fig. 3, which reflects all states for $\chi\leq 4.5\pi$. Guided by the minimum of potential $\omega$, as shown in Fig. 1, we conclude that for $\mu<0$, the phase-1 is favored over the metastable denser phase-2, whereas for $\mu>0$, the phase-2 becomes dominant. At the same time, Fig. 3 also shows unstable states $AB$ and $CD$ corresponding to local maxima of the potential $\omega$, when
@@
We stress that the simplifications adopted in the small-$\chi$ regime restrict the analysis to the unstable state only, thereby missing the complex – physically rich – picture of phenomena.
@@
We see that $c^{2}_{s}(a_{n})=0$ and $c^{2}_{s}(\chi)<0$ in unstable states, for example, for $\chi\in[0;\chi_{c}]$. Note that similar cases are known in the laboratory neg.sss-1 ; neg.sss-2 ; neg.sss-3 .
@@
Thomas–Fermi approximation. It involves disregarding the contribution of quantum fluctuations $\Delta_{\xi}\chi$ in Eq. (19), while considering the gravitational interaction, which leads to the set:
@@
It is easy to see that the spatial development stops when the first term vanishes due to $j_{1}(\chi)=0$, which happens at $\chi=a_{2}$. Such $\chi$ is actually reached at $\xi_{B}$ in Fig. 4. At $\xi>\xi_{B}$ we have to put $\chi=0$, but the gravitational potential behaves as $-1/\xi$.
@@ II.3 The excitation spectrum and instabilities
Given the quantum-fluid nature of BEC DM, let us analyze the spectrum of its small excitations. We reformulate Eq. (19) into the non-stationary Gross–Pitaevskii equation for the complex-valued wave function $X$, which depends on time $t$, as
@@
where the velocity vector ${\bf v}\equiv 2{\boldsymbol{\nabla}}\theta$ is expressed via the gradient operator ${\boldsymbol{\nabla}}$; and $\eta=\chi^{2}$ is the particle number density, as usual.
@@
This equation suggests that the background value $\eta_{0}$ evolves more slowly over time and space than the perturbation $\delta\eta$. In the absence of gravitation, it is possible to allow the homogeneity of matter, when $\eta_{0}$ is constant.
@@
where the wave vector squared $q^{2}=k^{2}/r_{0}^{2}$ and the mass density $\rho=\rho_{0}\eta_{0}$. The dimensional sound speed squared is
@@
Note that the obtained form of dispersion relation is well known H2019 ; Ch20 . And the models differ in the properties of the sound speed $v^{2}_{s}$.
@@
We immediately see that at the turning points in Fig. 3, when $\chi_{0}=a_{n}$ from the sequence $\{a_{n}\}$ in (32), the free-particle spectrum, $\omega_{k}^{(\pm)}=\pm k^{2}$, is recovered because $c^{2}_{s}(a_{n})=0$.
Since $j_{1}(\chi_{0})$ remains positive for $\chi_{0}\in(a_{2n-1};a_{2n})$, $n\in\mathbb{N}$, the excitation modes with $k^{2}<\chi_{0}j_{1}(\chi_{0})$ become unstable due to the imaginary $\omega_{k}^{(\pm)}$. These instability conditions are consistent with those predicted thermodynamically, in particular, for the segments $AB$ and $CD$ in Fig. 3. On the other hand, the physically constructive phase-1 and phase-2 correspond to $\chi_{0}$ in the intervals $(a_{2};a_{3})$ and $(a_{4};a_{5})$, where $j_{1}(\chi_{0})<0$ and $c^{2}_{s}>0$. Thereby, Eq. (65) also confirms the sustainability of these states.
@@
For a nonzero parameter $A\ll 1$, suppression of the gravitational Jeans instability in (62) necessitates the introduction of a lower bound $k_{J}$ on the wave numbers $k$. However, due to the substantially inhomogeneous DM distribution and the need for a nonperturbative approach, the limiting size of a stable DM cloud cannot be reliably determined based on $k_{J}$. Therefore, studies using time-dependent numerical analysis are needed to refine Jeans limit in size and mass. For this reason, we adopt the known limits for cold boson stars BGZ84 ; UTB02 and conservatively estimate the maximum stable mass as $M\sim 0.6\hbar c/(Gm)$.
@@ III.1 Solving the field equations and
Due to the nonlinearity of Eq. (19), the initial conditions of the spatial evolution of the field $\chi$ cannot be given arbitrarily, but must be determined. Requiring the absence of cuspidity at the origin, i.e. $\chi(0)<\infty$, $\chi^{\prime}(0)=0$, and $\chi^{\prime\prime}(0)<0$ for given $A$ and $\nu$, the finite initial value $\chi(0)=z$ should satisfy the equation:
@@
It is derived by substituting the approximation $\chi(\xi)=\chi(0)+\chi^{\prime\prime}(0)\xi^{2}/2$ for $\xi\to 0$ into (19)-(20) and by finding $\chi^{\prime\prime}(0)$ GN23 . Note that the condition $\chi^{\prime\prime}(0)<0$ is equivalent to requiring $\nu>j_{0}(z)$ and $\nu<\cos{z}$ for positive $\nu$.
Figure 5: Black curves are the solutions of Eq. (66) for $A=4.5\times 10^{-3}$. The red cross marks the initial condition for the blue curve in Fig. 4.
The form of Eq. (66) indicates that its solution $z$ exists for small $A$. For this reason, we take $A$ to be of the order $10^{-3}$. Physically, this means that gravitational interaction is significantly weaker than the self-interaction between particles. However, we immediately encounter the problem of the solution ambiguity as we see in Fig. 5. Moreover, the number of “islands”, as in Fig. 5, increases with decreasing $A$. As we shall see, this ambiguity leads to solutions with different thermodynamic properties.
@@
Before theoretically studying the effects in DM, we test the model ability to describe observables. It is natural to focus on the rotation curves (RCs) of galaxies Craciun2019 ; Kun2020 .
@@
Having obtained the optimal parameters $A$, $V_{0}$, and $r_{0}$, we can extract the dark axion mass $m$, the mass density scale $\rho_{0}$ and the central density $\rho_{c}=\rho_{0}\chi^{2}(0)$, the axion decay constant $f_{\text{a}}$ as
$m=\sqrt{\frac{A}{2}}\frac{\hbar}{V_{0}r_{0}},\quad\rho_{0}=\frac{V_{0}^{2}}{4\pi Gr_{0}^{2}},\quad f_{\text{a}}=\frac{m_{\text{Pl}}V_{0}^{2}}{\sqrt{2\pi A}},$ (69)
where $m_{\text{Pl}}\equiv\sqrt{\hbar c/G}=2.176434\times 10^{-8}$ kg.
Figure 6: Rotation curve for NGC 2366. Black dots corresponds the data from Oh08 . The solid curve is obtained on the base of the blue solution in Fig. 4.
@@
The boson mass $m$ allows us to roughly estimate the limiting mass of gravitationally stable DM sphere, $M_{c}\sim 0.6\hbar c/(Gm)\simeq 6.84\times 10^{13}M_{\odot}$, while the computed DM halo mass is about $10^{10}M_{\odot}<M_{c}$, where $M_{\odot}$ is the solar mass. This confirms the gravitational stability of DM in NGC 2366.
Another fitting of the RC of NGC 2366 in the BEC DM concept is shown in Harko2011 . There, the positive $s$-scattering length $a_{s}$ and wave number $k$, $k^{2}=Gm^{3}/(\hbar^{2}a_{s})$, are used. In contrast, if the pairwise interaction $\psi^{4}$ is naively isolated from potential (3) in the vicinity of $\psi=0$, which corresponds to an unstable state, we obtain a negative scattering length from the relation:
@@
On the other hand, by expanding potential (3) in a series in the vicinity of some $\psi$ belonging to a stable state, for example, phase-1, one may obtain a positive scattering length $a_{s}>0$, as it should be.
@@
The latter is involved in the Schrödinger-type equation $(-\Delta_{\xi}+W(\xi))\chi=\varepsilon\chi$, where $\varepsilon=1+2u$, and $\varepsilon\simeq-0.5867$ for the DM of NGC 2366.
Figure 8: Dependence of the grand potential density $\omega$ on order parameter $\chi$ for the blue solution in Fig. 4.
@@
Thereby, we have identified transient processes occurring in the tail of the DM distribution for NGC 2366. Further, we explore these processes in more detail, focusing on solutions $\chi(\xi)$ similar to the blue one shown in Fig. 4, and using a statistical approach.

### v2

@@ Abstract
The phase diagrams of ultralight dark matter (DM), modeled as a self-gravitating Bose–Einstein condensate with axionlike interaction, are studied. We classify stable, metastable, and unstable DM states over a wide range of condensate wave function amplitudes. It is shown that the axionlike interaction causes instability and an imaginary speed of sound at low amplitudes, whereas, in a specific high-amplitude band, DM attains a stable state capable of forming a dense solitonic core and suppressing quantum fluctuations in the surrounding galactic DM halo. These findings are corroborated by evaluating thermodynamic functions for DM in the dwarf galaxy NGC 2366 and its hypothetical analogs with different core-to-halo mass ratios. Distinct DM phase compositions respond differently to fluctuation-induced partial pressure, resulting in a first-order phase transition in a certain range of an interaction parameter. While the DM properties in NGC 2366 lie within the supercritical regime, the phase transition nonetheless provides a thermodynamic marker separating stable from unstable DM configurations. Once a dense core forms–reaching a threshold of about 12% of the total mass–the enhanced gravitation stabilizes the DM halo against fluctuations, while the internal pressure ensures core stability. In particular, we find that NGC 2366’s dense DM comprises roughly 19% of the DM mass while occupying only 4.7% of its total volume.
@@ I Introduction
One may observe parallels with the ensemble of oscillons Osc1 generated by the instanton self-interaction, leading in the absence of gravity to the three-dimensional sine-Gordon equation Man04 . Moreover, established results on axitons BZ19 ; KT94 ; ax5 are valuable for further comparison of the sets of states.
Neglecting the Universe’s evolution in describing current visible structures, we are thus dealing with the statistical properties of self-interacting DM (SIDM) within the model of self-gravitating Bose–Einstein condensate (BEC). Multifaceted phase diagrams should then reflect these properties, tracing transitions among distinct states and configurations.
Numerous studies have rigorously examined the properties of BEC DM with two- and three-particle interactions, using both exact numerical methods and various analytical approximations, such as the Thomas–Fermi approximation and the variational method (see Chavanis2 ; Peebles ; Bohmer ; Harko2011 ; Zang2018 ; Nazar ; SV24 ). The thermodynamic functions in these models have been carefully studied, and a new aspect of DM has emerged GKN20 ; GN21 ; GN23 : the prediction of a first-order phase transition between rarefied and dense states, which is driven by quantum fluctuation pressure and governed by coupling constants. Additionally, we would like to note the impact of the formation of composites from the initial DM particles GN23 ; GN22 ; GKKN , the quantum entanglement of parts of the DM halo GN21 ; Lee18 , and the emergence of superfluidity Berezhiani ; Ber25 .
@@
To constrain the parameters of our BEC DM model described by the Gross–Pitaevskii–Poisson equations, we consider the DM dominating the dwarf galaxy NGC 2366 as a reference. This nearby galaxy provides an opportunity to study reionization processes in vigorous low-metallicity starbursts, similar to those that drove cosmic reionization Hun01 ; Iz99 , and exhibits the DM halo profile inconsistent with the cuspy distributions predicted by cold DM models ($\Lambda$CDM) Oh08 . Earlier efforts have already addressed this core-cusp tension within the BEC DM model with pairwise interaction Harko2011 . In this regard, we intend to calibrate our model by fitting its rotation curve Oh08 ; Bl08 , and then generate a statistical ensemble of hypothetical DM distributions following from the model equations under similar conditions. By comparing different configurations using phase diagrams, we aim to reveal the uniqueness of the DM structure in NGC 2366, which forms the gravitational background for subgalactic processes. This analysis is expected to provide new insights into the galaxy’s DM and to test the diagnostic capabilities of the statistical approach.
The paper is structured as follows. In Sec. II we introduce the main equations of the model and, through various simplifications, analyze the properties essential for interpreting subsequent results. In Sec. III we explore nonperturbative solutions and their application to modeling the DM distribution in the dwarf galaxy NGC 2366. We then present our main results – statistical characteristics and phase diagrams based on the numerically obtained DM profiles. The final Sec. IV contains a summary of the findings and concluding remarks.
@@ II.1 The model equations
This prescription immediately leads to the dimensionless measure of gravitational interaction:
@@
With $v=\hbar^{3}c/(mf^{2}_{\text{a}})$ defined via the axion decay constant $f_{\mathrm{a}}$, it is then expected that in models featuring ultralight particles, $f_{\text{a}}$ may fall within the range $10^{18}\,\text{eV}<f_{\text{a}}<10^{21}\,\text{eV}$ Chavanis2 .
@@
In our study, $\nu$ is regarded as a free parameter, reflecting the arbitrariness of $\tilde{\mu}$.
@@ II.2 The particular cases
Parameters $\nu_{1}$-$\nu_{5}$ determine the local criticality of the curves in Fig. 1. Their values are established through conditions:
@@
Let us analyze the $p-\mu$ phase diagram in Fig. 3, which reflects all states for $|\chi|\leq 4.5\pi$. Guided by the minimum of potential $\omega$, as shown in Fig. 1, we conclude that for $\mu<0$, the phase-1 is favored over the metastable denser phase-2, whereas for $\mu>0$, the phase-2 becomes dominant. At the same time, Fig. 3 also shows unstable states $AB$ and $CD$ corresponding to local maxima of the potential $\omega$, when
@@
Thus, the simplifications adopted in the small-$\chi$ regime restrict the analysis to the unstable state only, thereby missing the complex – physically rich – picture of phenomena.
@@
We see that $c^{2}_{s}(a_{n})=0$, and $c^{2}_{s}(\chi)<0$ in unstable states, for example, for $\chi\in[0;\chi_{c}]$. Note that similar situations are known in the laboratory neg.sss-1 ; neg.sss-2 ; neg.sss-3 .
@@
Thomas–Fermi approximation. It involves disregarding the contribution of quantum fluctuations $\Delta_{\xi}\chi$ in Eq. (19), while considering the gravitational interaction, which leads to the set of equations:
@@
It is easy to see that the spatial development stops when the first term vanishes due to $j_{1}(\chi)=0$, which happens at $\chi=a_{2}$. Such $\chi$ is actually reached at $\xi_{B}$ in Fig. 4. At $\xi>\xi_{B}$ we have to put $\chi=0$, but the gravitational potential behaves as $1/\xi$.
@@ II.3 The excitation spectrum and instabilities
Given the quantum-fluid nature of BEC DM, let us analyze the spectrum of its small excitations. We reformulate Eq. (19) into the non-stationary Gross–Pitaevskii equation for the complex-valued wave function $X$, which depends on dimensionless time $t$, as
@@
where the velocity vector ${\bf v}\equiv 2{\boldsymbol{\nabla}}\theta$ is expressed via the gradient operator ${\boldsymbol{\nabla}}$; and $\eta=\chi^{2}$ is the particle number density, as before.
@@
This equation admits that the background value $\eta_{0}$ evolves more slowly over time and space than the perturbation $\delta\eta$. In the absence of gravitation, the homogeneity of matter is allowed, that is, when $\eta_{0}$ is constant.
@@
where the wave vector squared $q^{2}=k^{2}/r_{0}^{2}$, the mass density $\rho=\rho_{0}\eta_{0}$, and the squared sound speed
@@
This dispersion relation is well known Ber25 ; H2019 ; Ch20 . And the models differ in the properties of the sound speed $v_{s}$.
@@
We immediately see that at the turning points in Fig. 3, i.e. $\chi_{0}=a_{n}$ from the sequence $\{a_{n}\}$ in (32), the sound speed vanishes, $c^{2}_{s}(a_{n})=0$, and the free-particle spectrum, $\omega_{k}^{(\pm)}=\pm k^{2}$, is recovered.
Since $j_{1}(\chi_{0})$ remains positive for $\chi_{0}\in(a_{2n-1};a_{2n})$, $n\in\mathbb{N}$, the excitation modes with $k^{2}<\chi_{0}j_{1}(\chi_{0})$ become unstable due to the imaginary $\omega_{k}^{(\pm)}$. These instability conditions are consistent with those predicted thermodynamically, in particular, for the segments $AB$ and $CD$ in Fig. 3. On the other hand, the physically constructive phase-1 and phase-2 correspond to $\chi_{0}$ in the intervals $(a_{2};a_{3})$ and $(a_{4};a_{5})$, where $j_{1}(\chi_{0})<0$ and $c^{2}_{s}>0$. Thereby, Eq. (65) also confirms the sustainability of these states. In this case, the Bogolyubov spectrum is obtained, similar to that in superfluidity models Ber25 .
@@
For a nonzero parameter $A\ll 1$, to overcome the gravitational Jeans instability at $k\to 0$ in (62), it is necessary to introduce a lower bound $k_{J}$ for wave numbers $k$. However, due to the substantially inhomogeneous DM distribution and the need for a nonperturbative approach, the limiting size of a stable DM cloud cannot be reliably determined based on $k_{J}$. Therefore, studies using time-dependent numerical analysis are needed to refine Jeans limit in size and mass. For this reason, we adopt the known limits for cold boson stars BGZ84 ; UTB02 and conservatively estimate the maximum stable mass as $M\sim 0.6\hbar c/(Gm)$.
@@ III.1 Solving the field equations and
Due to the nonlinearity of Eq. (19), the initial conditions of the spatial evolution of the field $\chi$ cannot be given arbitrarily, but must be determined. By requiring the absence of cuspidity at the origin, i.e. $\chi(0)<\infty$, $\chi^{\prime}(0)=0$, and $\chi^{\prime\prime}(0)<0$ for given $A$ and $\nu$, the finite initial value $\chi(0)=z$ should satisfy the equation:
@@
This is derived by substituting the expansion $\chi(\xi)=\chi(0)+\chi^{\prime\prime}(0)\xi^{2}/2$ for $\xi\to 0$ into (19)-(20) and by finding $\chi^{\prime\prime}(0)$ GN23 . Note that the condition $\chi^{\prime\prime}(0)<0$ is equivalent to requiring $j_{0}(z)<\nu<\cos{z}$ for positive $\nu$.
Figure 5: Black curves represent the solutions of Eq. (66) for $A=4.5\times 10^{-3}$. The red cross marks the initial condition for the blue curve in Fig. 4.
The form of Eq. (66) indicates that its solution $z$ exists for small $A$. For this reason, we take $A$ to be of the order of $10^{-3}$. Physically, this means that gravitational interaction is significantly weaker than the self-interaction between particles. However, we immediately encounter the problem of the solution ambiguity as we see in Fig. 5. Moreover, the number of “islands”, as in Fig. 5, increases with decreasing $A$. As we shall see, this ambiguity leads to solutions with different thermodynamic properties.
@@
Before theoretically studying the effects in DM, we test the model ability to describe observables. It is natural to focus on the rotation curves (RCs) of galaxies. BEC DM models have already been successfully employed to reproduce RCs for galaxies from the SPARK database Harko2011 ; Craciun2019 ; Kun2020 .
@@
Having obtained the optimal parameters $A$, $V_{0}$, and $r_{0}$, we can extract the dark axion mass $m$, the mass density scale $\rho_{0}$ defining the central density $\rho_{c}=\rho_{0}\chi^{2}(0)$, and the axion decay constant $f_{\text{a}}$ as
$m=\sqrt{\frac{A}{2}}\frac{\hbar}{V_{0}r_{0}},\quad\rho_{0}=\frac{V_{0}^{2}}{4\pi Gr_{0}^{2}},\quad f_{\text{a}}=\frac{m_{\text{P}}V_{0}^{2}}{\sqrt{2\pi A}},$ (69)
where $m_{\text{P}}\equiv\sqrt{\hbar c/G}=2.176434\times 10^{-8}$ kg.
To facilitate estimation, we provide the formulas:
$m$ $=$ $0.4287\times 10^{-22}~{}\text{eV}~{}c^{-2}~{}\left[\frac{A}{10^{-3}}\right]^{1/2}$
×$\times\left[\frac{r_{0}}{1~{}\text{kpc}}\right]^{-1}\left[\frac{V_{0}}{1~{}\text{km}/\text{s}}\right]^{-1},$
$\rho_{0}$ $=$ $0.1252\times 10^{-23}~{}\text{kg}~{}\text{m}^{-3}~{}\left[\frac{r_{0}}{1~{}\text{kpc}}\right]^{-2}\left[\frac{V_{0}}{1~{}\text{km}/\text{s}}\right]^{2},$
$f_{\text{a}}$ $=$ $0.1714\times 10^{19}~{}\text{eV}~{}\left[\frac{A}{10^{-3}}\right]^{-1/2}\left[\frac{V_{0}}{1~{}\text{km}/\text{s}}\right]^{2}.$
Figure 6: Rotation curve for NGC 2366. Black dots corresponds to the data from Oh08 . The solid curve follows from the blue solution in Fig. 4.
@@
The boson mass $m$ allows us to roughly estimate the limiting mass of gravitationally stable DM sphere, $M_{c}\sim 0.6\hbar c/(Gm)\simeq 6.84\times 10^{13}M_{\odot}$, while the computed DM halo mass is about $10^{10}M_{\odot}<M_{c}$, where $M_{\odot}$ is the solar mass. This indicates the gravitational stability of the DM in NGC 2366.
Another fitting of the RC of NGC 2366 in the BEC DM concept is shown in Harko2011 . There, the positive $s$-scattering length $a_{s}$ and wave number $k$, $k^{2}=Gm^{3}/(\hbar^{2}a_{s})$, are used. In contrast, if the pairwise interaction $\psi^{4}$ is naively isolated from potential (3) in the vicinity of $\psi=0$, which corresponds to an unstable state, we obtain a negative, usually underestimated, scattering length from the relation:
@@
On the other hand, by expanding potential (3) in a series in the vicinity of some $\psi$ belonging to a stable state, for example, phase-1, one may obtain a positive scattering length $a_{s}>0$, as it is expected.
@@
The latter is involved in the Schrödinger-type equation $[-\Delta_{\xi}+W(\xi)]\chi=\varepsilon\chi$, where $\varepsilon=1+2u$, and $\varepsilon\simeq-0.5867$ for the DM in NGC 2366.
Figure 8: Dependence of the grand potential density $\omega$ on the order parameter $\chi$ determined by the blue solution in Fig. 4.
@@
Thereby, we have identified transient processes occurring in the tail of the DM distribution in NGC 2366. Further, we explore these processes in more detail, focusing on solutions $\chi(\xi)$ similar to the blue one shown in Fig. 4, and using a statistical approach.